Session: Feb 2005 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Fri Feb 25 07:40:11 2005) Success :
All the equipment was good, however starting from about 02UT was snowing. The dish had a ligth layer of snow and Tsys was about two times as usual.

Torun (K. Borkowski at Fri Feb 25 08:41:45 2005) Success : RECO HIGH
First tape not very smoothly packed. Snowing during experiment (higher Tsys). No other known problems.

Hartebeesthoek (Jonathan Quick at Fri Feb 25 08:54:11 2005) Success :
No problems detected.

Onsala (Michael Lindqvist at Fri Feb 25 10:22:43 2005) Success :
No known problems.

Arecibo (T. Ghosh and E. Momjian at Mon Feb 28 11:28:52 2005) Success :
No problems as far as we can see.

Effelsberg (A. Kraus at Tue Mar 1 16:05:25 2005) Success :

Jodrell_Bank (Alastair Gunn at Mon Mar 7 11:49:25 2005) Success :
MkII was used. No known problems.